OpenStudy (anonymous):

Write an equation in slope-intercept form for the line that satisfies the following condition. slope 4, and passes through (2, 20)

OpenStudy (anonymous):

\[y-y_1=m(x-x_1)\] with \(m=4,x_1=2,y_1=20\) gives \[y-20=4(x-2)\] you can solve for \(y\) if you choose

OpenStudy (anonymous):

will you solve for y so i can see please

OpenStudy (anonymous):

\[y-20=4(x-2)\] use the distributive property to remove the parentheses \[y-20=4x-8\] add 20 to both sides \[y=4x+12\]

OpenStudy (anonymous):

thank you i am new to algebra

OpenStudy (anonymous):

yw